A boy on board a cruise ship drops a 30.0 gm marble into the ocean. If the

resistive force proportionality constant is 0.500 kg/s, what is the terminal speed

of the marble in m/s?


1
Expert's answer
2021-02-24T16:58:33-0500

"F=mg,\\ v=\frac Fk=\frac{mg}{k},\\
v=\frac{0.03\cdot 9.81}{0.5}=0.589~\frac{\text{m}}{\text{s}}."